# Controlled neuromodulation naturally interacts with homeostasis to produce robust and modulable neuronal function

### v2

@@ Abstract
Neurons rely on two interdependent mechanisms – homeostasis and neuromodulation – to maintain robust and adaptable functionality. Homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals carried by neuromodulators. Combining these mechanisms in conductance-based models often produces unreliable outcomes, particularly when sharp neuromodulation interferes with homeostatic tuning. This study explores how a biologically inspired neuromodulation controller can harmonize with homeostasis to ensure reliable neuronal function. Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while homeostasis simultaneously maintains target intracellular calcium levels. Unlike sharp neuromodulation, the neuromodulation controller integrates activity-dependent feedback through mechanisms mimicking G-protein-coupled receptor cascades. The interaction between these controllers critically depends on the existence of an intersection in conductance space, representing a balance between target calcium levels and neuromodulated firing patterns. Maximizing neuronal degeneracy enhances the likelihood of such intersections, enabling robust modulation and compensation for channel blockades. We further show that this controller pairing extends to network-level activity, reliably modulating central pattern generators rhythmic activity. This study highlights the complementary roles of homeostasis and neuromodulation, proposing a unified control framework for maintaining robust and adaptive neural activity under physiological and pathological conditions.
@@
Neurons must maintain stable activity while adapting to changing demands, relying on two mechanisms: homeostasis, which stabilizes activity, and neuromodulation, which adjusts behavior to external signals. These mechanisms often interact, but their improper coordination can lead to dysfunction. This study uses computational models to show how controlled neuromodulation – mimicking biological feedback – can be harmonized with controlled homeostasis to ensure robust neuronal function. The system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks even when faced with neuronal parameter variability. These insights not only deepen our understanding of neural robustness, but also suggest safer pharmacological strategies targeting neuromodulatory pathways, offering potential breakthroughs in treating neurological disorders without disrupting essential cellular mechanisms.
@@ Introduction
Neuromodulation coexists with homeostatic plasticity, a process that gradually adjusts neuronal membrane properties to maintain a target activity level [8, 29, 37, 7, 50, 49, 51]. Neuronal cells are dependent on homeostatic mechanisms to maintain optimal functionality throughout the extended useful life of mammals [32]. Despite continuous turnover of transmembrane proteins, such as ion channels and receptors, occurring on varying temporal scales ranging from hours to weeks [39], the robustness and excitability of neurons must remain undisturbed [31]. Among homeostatic mechanisms, intrinsic homeostatic plasticity adjusts ion channel conductances in response to intracellular calcium levels [36, 41]. However, since both neuromodulation and homeostasis act on the same targets (ion channels) and may have competing objectives, understanding their interaction is essential [30, 31].
@@
We propose to study the interaction between neuromodulation and homeostasis using conductance-based models. These models provide a biophysically accurate representation of neuronal dynamics by describing the electrical properties of the neuronal membrane. The membrane is modeled as an RC circuit, while ion channels are represented as nonlinear conductances that evolve according to experimentally derived activation and inactivation kinetics [22]. Conductance-based models remain the most faithful mathematical representations of biological neurons, as they allow for the direct incorporation of ion channel dynamics extracted from experimental data, following the methodology first introduced by Hodgkin and Huxley [19]. Over the years, numerous conductance-based models have been developed to describe a variety of neuronal cell types across different species and brain regions [25, 43, 40, 52, 12, 23, 34]. This study focuses on a conductance-based model of stomatogastric ganglion (STG) neurons [25]. STG neurons are key components of rhythmic circuits that generate motor patterns driving stomach muscle contractions in crustaceans. This model was chosen for its ability to reproduce diverse firing patterns – ranging from tonic spiking to bursting – and its rich set of ion channels, which enable degeneracy. Specifically, the model includes classical fast sodium and delayed rectifier potassium channels, A-type and calcium-activated potassium channels, T-type and slow calcium channels, as well as hyperpolarization-activated H-channels.
@@
We model intrinsic homeostasis by using the homeostatic controller proposed by [38], which adjusts all ionic conductances up or down based on the neuron average intracellular calcium level. If the calcium level is lower than a target reference value, the controller increases ionic conductances according to homeostatic tuning rules, and conversely, decreases them if the calcium level is above the target. Importantly, these rules ensure that conductances are tuned along a single axis, preserving their relative ratios and maintaining correlations between them. This self-tuning mechanism allows neurons to counteract perturbations while preserving stable activity.
@@
Neuromodulation is modeled using two different approaches. On the one hand, we model neuromodulation as an acute change in target conductance values leading to a desired neuromodulated state in degenerate populations, which we call ”sharp neuromodulation”. On the other hand, we implement the neuromodulation controller developed by [14], which is designed to enable robust modulation of neuronal activity. We call this approach ”controlled neuromodulation”. One of its key properties is that it accounts for neuronal degeneracy, meaning that neurons with different ion channel compositions but similar activity patterns respond in a consistent manner to neuromodulation [15]. This feature is particularly important because biological neuronal populations exhibit significant variability in their underlying conductance parameters while maintaining functional stability. By leveraging degeneracy, the neuromodulation controller can apply the same control signal across a population of neurons with diverse conductance profiles, ensuring reliable and predictable neuromodulation.
@@
We first investigate the interaction between intrinsic homeostatic plasticity and neuromodulation at the single-neuron level and show that combining intrinsic homeostasis with either sharp or controlled neuromodulation leads to vastly different outcomes. Secondly, we provide a mechanistic explanation for these different outcomes. We then extend our analysis to rhythmic neuronal networks, where neuromodulation plays a critical role in adjusting oscillatory patterns. By carefully integrating homeostatic and neuromodulatory control mechanisms, we aim to develop a unified framework that combines the strengths of both approaches. This integrated controller enables robust modulation of neuronal activity while ensuring long-term stability, shedding light on how neuromodulation and homeostasis coexist in biological systems.
@@ Results
Interference between homeostasis and sharp neuromodulation leads to pathological outcomes
We first examine the interplay between a state-of-the-art model of cellular homeostasis [38] and sharp neuromodulation [13]. Here, sharp neuromodulation is defined as a direct, externally driven modification of a subset of ion channel maximal conductances, resulting in a change in neuronal spiking behavior. We specifically focus on a robust transition from tonic spiking to bursting in degenerate neuronal populations, i.e. populations composed of neurons that exhibit similar firing behaviors despite differences in underlying conductance profiles.
To explore this interaction, we applied sharp neuromodulation to a population of degenerate neurons generated using the dataset from [13] (Fig 1). Sharp neuromodulation consisted of directly modifying the maximal conductances of A-type potassium ($\bar{g}_{\mathrm{A}}$) and slow calcium ($\bar{g}_{\mathrm{CaS}}$) channels to transition neurons from tonic spiking to bursting. We then allowed homeostatic compensation to unfold. Since bursting increases overall neuronal activity, intracellular calcium levels rise, triggering a decrease in all conductances according to homeostatic tuning rules. Three representative conductance trajectories ($N=200$) are shown in Fig 1 (top), while Fig 1 (bottom) illustrates calcium dynamics across the population (note that to facilitate observation of both mechanisms within a single trace, the homeostatic controller was accelerated compared to its original implementation). Despite initially similar firing patterns and a similar outcome of sharp neuromodulation, homeostatic compensation produced highly variable and often unreliable outcomes in terms of modulated activity, while still regulating intracellular calcium. Depending on the initial conductance state, neuromodulation led to new states where homeostasis could either maintain functional activity, result in pathological behavior, or drive neurons into unphysiological regimes. This finding underscores a fundamental property of degeneracy: neurons with different underlying conductances can exhibit similar activity yet respond differently to perturbations, such as neuromodulation.
Fig 1: Interference between homeostasis and sharp neuromodulation leads to pathological outcomes. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of 200 models (top). The corresponding mean intracellular calcium concentration over time (bottom) shows the target value (red dash-dotted line) is maintained both before and after sharp neuromodulation.
To understand the source of this unreliability, it is necessary to examine both mechanisms in more detail. The homeostatic model [38] relies on intracellular calcium as a feedback signal. The neuron continuously averages its intracellular calcium level over time and adjusts conductances accordingly: if calcium falls below a target level (red dash-dotted line in Fig 1 bottom), ion channel transcription and translation increase, raising all conductance values, and vice versa. This phenomenon happens on a relatively slow timescale, which can last up to days [36]. In contrast, sharp neuromodulation, implemented using the method from [13], directly modifies a subset of ion channel conductances, which upsets the homeostasis setpoint, defined by the intracellular calcium target level. As homeostasis kicks in to recover its calcium set point, there are no guarantees that the neuromodulated function is maintained as ion channel conductances are modified, which often leads to pathological or unphysiological outcomes.
This type of interference was anticipated in [38], where the authors demonstrated that homeostatic regulation can produce diverse responses to perturbations such as ion channel deletions. Depending on the initial conditions, a deletion may either preserve function, disrupt it, or lead to pathological compensation. Our results extend this idea by showing that even physiological disturbances, such as neuromodulation, can induce similar unpredictability when coupled with homeostatic regulation.
Controlled neuromodulation and homeostasis naturally lead to robust and modulable neuronal function
In biological systems, neuromodulation occurs through an intracellular cascade that indirectly links changes in neuromodulation concentration and changes in targeted conductance values using an intermediate signal such as e.g. cyclic amp. We have recently shown that this indirect action of neuromodulation on conductance values could be viewed as an intracellular controller (Fig 2A, top), which largely improves reliability of neuromodulatory action on highly degenerate populations [14]. We call this mechanism ”controlled neuromodulation” to contrast with the sharp neuromodulation described above. Unlike sharp neuromodulation, which imposes instantaneous conductance changes, the neuromodulation controller provides continuous and adaptive modulation, ensuring stable functional transitions.
Fig 2: Controlled neuromodulation and homeostasis naturally lead to robust and modulable neuronal function.
@@
A. Schematic representation of neuromodulation (left) and homeostasis (right) cascades. Neuromodulators bind to G-protein receptors, triggering complex signaling pathways that selectively affect subsets of ion channels. In contrast, homeostasis senses intracellular calcium levels and regulates them by adjusting all conductances uniformly.
B. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during homeostasis with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$) for a degenerate population of 200 models (left). The corresponding mean intracellular calcium concentration over time (right) shows the target value (red dash-dotted line) is maintained both before and after controlled neuromodulation.
@@
Using the same protocol as for sharp neuromodulation, we now study the interaction between controlled neuromodulation and homeostasis. Fig 2B illustrates the transition from tonic spiking to bursting in a degenerate population of STG neuron models, identical to those in Fig 1, but when using controlled neuromodulation instead of sharp neuromodulation. Initially, all neurons exhibit tonic firing, with homeostasis tuning conductances along a tonic spiking homogeneous scaling line to maintain a target calcium level. Midway through the simulation, controlled neuromodulation is applied, gradually adjusting $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ to induce bursting. As in the case of sharp neuromodulation, this transition increases intracellular calcium levels, leading to an adaptive shift in conductance values. Homeostatic regulation then readjusts overall conductance magnitudes, but here the interaction between homeostasis and controller neuromodulation ensures that the newly established neuromodulated state is maintained. Ultimately, the population reaches a stable configuration where both the neuromodulated function and intracellular calcium homeostasis are fulfilled.
To demonstrate the generality of this approach, we applied the same framework to a midbrain dopaminergic neuron model adapted from [43] (see Supporting Information, S1 Appendix). As in the STG model, a transition from tonic spiking (pacemaking) to strong bursting was induced by modulating N-type and L-type calcium conductances ($\bar{g}_{\mathrm{CaN}}$ and $\bar{g}_{\mathrm{CaL}}$) through controlled neuromodulation. Homeostatic compensation continuously regulated calcium levels, ensuring a reliable and robust outcome despite initial conductance degeneracy within the neuronal population.
@@
Controlled neuromodulation tunes the direction of homeostasis-driven changes in conductance values to ensure stable neuromodulated function during homeostatic regulation
@@
To study the mechanisms underlying the constructive interaction between controlled neuromodulation and homeostasis, we plot the conductance trajectories of the simulations from Fig 1 in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking homeostatic compensation for both sharp and controlled neuromodulation (Fig 3A). Initially, as calcium levels remain below the threshold, homeostasis increases the modulated conductances along a homogeneous scaling direction characteristic of tonic spiking. This direction corresponds to a line intersecting the origin, with variability due to conductance degeneracy within the population, and its slope is determined by the relative time constants of mRNA transcription in the homeostasis model, with steady-state analysis predicting $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$ [38]. When neuromodulation is applied, all conductances are adjusted to reshape their ratios and induce bursting (Fig 3B). Importantly, neuromodulatory actions follow the same direction across neurons but vary in magnitude depending on individual neuronal states [13]. This modulation leads to a sharp increase in intracellular calcium levels. Once neuromodulation ceases, the long-term outcome depends on the type of neuromodulation applied (Fig 3C).
Fig 3: Controlled neuromodulation tunes the direction of homeostasis-driven changes in conductance values to ensure stable neuromodulated function during homeostatic regulation.
@@
A. Trajectories of the STG population from Fig 1 in the modulated conductance space during homogeneous scaling in tonic spiking only. As expected from homeostasis, all neurons move in the same direction. White circles represent the initial conditions at the start of the step, and the green triangle marks the final conditions.
B. Same as panel A, but during the brief period when neuromodulation is active. All trajectories are parallel to one another.
C. Same as panels A and B, but during the homeostasis phase following neuromodulation. The outcome depends on the type of neuromodulation: sharp (left) or controlled (right). Sharp neuromodulation results in non-robust outcomes, whereas controlled neuromodulation preserves robust neural function.

### v3

# Activity-dependent neuromodulation and calcium homeostasis cooperate to produce robust and modulable neuronal function

@@ Abstract
Neurons rely on two interdependent mechanisms — homeostasis and neuromodulation — to maintain robust and adaptable functionality. Calcium homeostasis stabilizes neuronal activity by adjusting ionic conductances, whereas neuromodulation dynamically modifies ionic properties in response to external signals carried by neuromodulators. Combining these mechanisms in conductance-based models often produces unreliable outcomes, particularly when sharp neuromodulation interferes with calcium-homeostatic tuning. This study explores how a biologically inspired neuromodulation controller can harmonize with calcium homeostasis to ensure reliable neuronal function. Using computational models of stomatogastric ganglion and dopaminergic neurons, we demonstrate that controlled neuromodulation preserves neuronal firing patterns while calcium homeostasis simultaneously maintains target intracellular calcium levels. Unlike sharp neuromodulation, the neuromodulation controller integrates activity-dependent feedback through mechanisms mimicking G-protein-coupled receptor cascades. The interaction between these controllers critically depends on the existence of an intersection in conductance space, representing a balance between target calcium levels and neuromodulated firing patterns. Maximizing neuronal degeneracy enhances the likelihood of such intersections, enabling robust modulation and compensation for channel blockades. We further show that this controller pairing extends to network-level activity, reliably modulating the rhythmic activity of central pattern generators. This study highlights the complementary roles of calcium homeostasis and neuromodulation, proposing a unified control framework for maintaining robust and adaptive neural activity under physiological and pathological conditions.
@@
Neurons must maintain stable activity while adapting to changing demands, relying on two mechanisms: calcium homeostasis, which stabilizes activity, and neuromodulation, which adjusts behavior to external signals. These mechanisms often interact, but their improper coordination can lead to dysfunction. This study uses computational models to show how controlled neuromodulation — mimicking biological feedback — can be harmonized with calcium homeostasis to ensure robust neuronal function. The system compensates for channel blockades and maintains critical activity patterns for both single neurons and networks despite neuronal parameter variability. These insights not only deepen our understanding of neural robustness, but also suggest safer pharmacological strategies targeting neuromodulatory pathways, offering potential breakthroughs in treating neurological disorders without disrupting essential cellular mechanisms.
@@ Introduction
Neuromodulation coexists with homeostatic plasticity, a process that gradually adjusts neuronal membrane properties to maintain a target activity level [9, 36, 44, 8, 62, 61, 63]. Neuronal cells are dependent on homeostatic mechanisms to maintain optimal functionality throughout the extended useful life of mammals [39]. Despite continuous turnover of transmembrane proteins, such as ion channels and receptors, occurring on varying temporal scales ranging from hours to weeks [46], the robustness and excitability of neurons must remain undisturbed [38]. Among homeostatic mechanisms, intrinsic homeostatic plasticity adjusts ion channel conductances in response to intracellular calcium levels [43, 49]. However, since both neuromodulation and calcium homeostasis act on the same targets (ion channels) and may have competing objectives, understanding their interaction is essential [37, 38].
@@
We propose to study the interaction between neuromodulation and calcium homeostasis using conductance-based models. These models provide a biophysically accurate representation of neuronal dynamics by describing the electrical properties of the neuronal membrane. The membrane is modeled as an RC circuit, while ion channels are represented as nonlinear conductances that evolve according to experimentally derived activation and inactivation kinetics [28]. Conductance-based models remain the most faithful mathematical representations of biological neurons, as they allow direct incorporation of ion channel dynamics extracted from experimental data, following the methodology first introduced by Hodgkin and Huxley [24]. Over the years, numerous conductance-based models have been developed to describe a variety of neuronal cell types across different species and brain regions [31, 52, 47, 64, 13, 29, 41]. This study focuses on a conductance-based model of stomatogastric ganglion (STG) neurons [31]. STG neurons are key components of rhythmic circuits that generate motor patterns driving stomach muscle contractions in crustaceans. This model was chosen for its ability to reproduce diverse firing patterns — ranging from tonic spiking to bursting — and its rich set of ion channels, which enable degeneracy. Specifically, the model includes classical fast sodium and delayed rectifier potassium channels, A-type and calcium-activated potassium channels, T-type and slow calcium channels, as well as hyperpolarization-activated H-channels.
@@
We model intrinsic homeostasis using the calcium-homeostatic controller proposed by [45], which adjusts all ionic conductances based on the deviation of the neuron average intracellular calcium level from a target set-point. We refer to this mechanism as calcium homeostasis throughout this work, to distinguish it from other forms of homeostatic plasticity. Specifically, if the calcium level is lower than a target reference value, the controller increases ionic conductances according to calcium-homeostatic tuning rules, and conversely, decreases them if the calcium level is above the target. Importantly, these rules ensure that conductances are tuned along a single axis, preserving their relative ratios and maintaining correlations between them. This self-tuning mechanism allows neurons to counteract perturbations while preserving stable activity.
@@
Neuromodulation is modeled using two different approaches. On the one hand, we model neuromodulation as a hardwired change in target conductance values leading to a desired neuromodulated state in degenerate populations, which we call “sharp neuromodulation”. On the other hand, we implement the neuromodulation controller developed by [16], which is designed to enable robust modulation of neuronal activity. We call this approach “controlled neuromodulation”. One of its key properties is that it accounts for neuronal degeneracy, meaning that neurons with different ion channel compositions but similar activity patterns respond in a consistent manner to neuromodulation [17]. This feature is particularly important because biological neuronal populations exhibit significant variability in their underlying conductance parameters while maintaining functional stability. By leveraging degeneracy, the neuromodulation controller can apply the same control signal across a population of neurons with diverse conductance profiles, ensuring reliable and predictable neuromodulation.
@@
We first describe the calcium-homeostatic controller and the two neuromodulation approaches, sharp and controlled, highlighting their distinct dependence on neuronal activity. We then investigate their interaction at the single-neuron level and show that combining intrinsic homeostasis with either sharp or controlled neuromodulation leads to vastly different outcomes. Next, we provide a mechanistic explanation for these different outcomes. We then extend our analysis to rhythmic neuronal networks, where neuromodulation plays a critical role in adjusting oscillatory patterns. By carefully integrating homeostatic and neuromodulatory control mechanisms, we aim to develop a unified framework that combines the strengths of both approaches. This integrated controller enables robust modulation of neuronal activity while ensuring long-term stability, shedding light on how neuromodulation and calcium homeostasis coexist in biological systems.
@@ Results
Calcium homeostasis and neuromodulation: two complementary regulatory mechanisms
We begin by describing the two regulatory mechanisms that are central to this study: calcium homeostasis and neuromodulation. Both mechanisms act on ion channel conductances, but they differ fundamentally in their objectives, timescales, and dependence on neuronal activity.
The calcium-homeostatic model [45] relies on intracellular calcium as a feedback signal, as illustrated in Fig 1A. The neuron continuously averages its intracellular calcium level over time and adjusts conductances accordingly: if calcium falls below a target level, ion channel transcription and translation increase, raising all conductance values, and vice versa. This phenomenon happens on a relatively slow timescale, which can last up to days [43]. Importantly, the calcium-homeostatic tuning rules ensure that conductances are tuned along a single axis, preserving their relative ratios and maintaining correlations between them. Starting from any initial condition, this controller drives the cell along a line in conductance space known as the homogeneous scaling line. This self-tuning mechanism allows neurons to counteract perturbations while preserving stable activity.
@@
Fig 1: Schematic representations of calcium homeostasis, sharp and controlled neuromodulation. A. Schematic representation of calcium-homeostatic cascade. It senses intracellular calcium levels and regulates them by adjusting all conductances uniformly. B. Schematic representations of sharp (left) and controlled (right) neuromodulation. In sharp neuromodulation, neuromodulators bind to G-protein-coupled receptors, which directly affect subsets of ion channels independently of neuronal activity. In contrast, during controlled neuromodulation, G-protein-coupled receptors trigger complex signaling pathways that selectively affect subsets of ion channels in an activity-dependent way, hence producing a separate feedback loop over neuronal activity, acting in parallel with the calcium-homeostatic loop. Created in Bio Render. Fyon, A. (2026) https://BioRender.com/6q3irjy.
In contrast, neuromodulation modifies a subset of ion channel conductances to reshape firing patterns. We consider two distinct implementations: sharp neuromodulation and controlled neuromodulation. Sharp neuromodulation is defined as a direct, hardwired, externally driven modification of a subset of ion channel maximal conductances, applied independently of neuronal activity (Fig 1B left). In this approach, neuromodulators bind to G-protein-coupled receptors, which directly affect subsets of ion channels without incorporating feedback from the neuron current activity state. Sharp neuromodulation, implemented using the method from [15], thus upsets the homeostasis setpoint, defined by the intracellular calcium target level. As calcium homeostasis subsequently acts to recover its calcium set point, there are no guarantees that the neuromodulated function is maintained as ion channel conductances are modified.
Alternatively, we can take into account that, in biological systems, neuromodulation occurs through an intracellular cascade that indirectly links changes in neuromodulator concentration and changes in targeted conductance values using intermediate signals such as, e.g., cyclic adenosine monophosphate (cAMP) or intracellular calcium. We have recently shown that this indirect action of neuromodulation on conductance values could be viewed as an intracellular controller (Fig 1B right), which greatly improves the reliability of neuromodulatory action on highly degenerate populations [16]. We call this mechanism “controlled neuromodulation” to contrast with the sharp neuromodulation described above. Unlike sharp neuromodulation, which imposes activity-independent conductance changes, the neuromodulation controller provides continuous and adaptive modulation in an activity-dependent manner, ensuring stable functional transitions. This approach is biologically inspired, as there is experimental evidence that neuromodulation acts in an activity-dependent manner [30, 54, 19, 18, 56, 37]. Specifically, the neuromodulation controller models the role of G-protein-coupled receptors and second messengers in a non-specific way, embedding feedback within the metabolic neuromodulation cascade [33, 55]. During controlled neuromodulation, G-protein-coupled receptors trigger complex signaling pathways that selectively affect subsets of ion channels in an activity-dependent way, hence producing a separate feedback loop over neuronal activity, acting in parallel with the calcium-homeostatic loop.
At its core, the controller exploits Dynamic Input Conductances (DICs), a low-dimensional representation that links ion channel conductances to firing patterns (see Materials and Methods for details). DICs reduce the high-dimensional space of ion channel conductances to three voltage-dependent feedback gains, each associated with a distinct timescale (fast, slow, and ultraslow), whose values at threshold voltage reliably determine the firing pattern [11]. The controller operates in two stages. First, a feedforward step uses this linear mapping to compute neuron-specific reference values for the modulated conductances, given the desired firing pattern and the current state of all other (unmodulated) conductances. Second, a feedback step in the form of a Proportional-Integral (PI) controller tracks these references by regulating ion channel expression. Because the feedforward step depends on the full conductance state of the neuron, the controller naturally adapts its output to each individual neuron within a degenerate population, producing different conductance adjustments that all converge to the same target activity.
Importantly, the controller is agnostic to the specific neuromodulator identity and intracellular signaling cascade involved. This abstraction is justified by experimental evidence showing that diverse signaling pathways converge onto the same functional targets [37]. For instance, in the STG, multiple neuromodulators acting through distinct G-protein-coupled receptors and second messenger pathways (involving cAMP or calcium-dependent cascades) all converge onto the same modulator-activated inward current $I_{\mathrm{MI}}$ [60]. In this context, the feedforward step captures the dependence of neuromodulatory effects on the current activity state of the neuron [30], while the feedback step (PI controller) models the temporal integration inherent to intracellular signaling cascades, where second messenger concentrations gradually adjust in response to sustained receptor activation. Rather than modeling any particular cascade, our controller captures the functional principle common to these diverse pathways: an activity-dependent feedback loop that senses neuronal output and adjusts a targeted subset of conductances to reach a desired activity pattern.
Sharp neuromodulation interferes with calcium homeostasis whereas controlled neuromodulation naturally leads to robust and modulable neuronal function
Having established the two neuromodulation approaches, we now examine their interaction with calcium homeostasis. We specifically focus on a robust transition from tonic spiking to bursting in degenerate neuronal populations, i.e. populations composed of neurons that exhibit similar firing behaviors despite differences in underlying conductance profiles. We applied both types of neuromodulation to a population of degenerate neurons generated using the dataset from [15], modifying the maximal conductances of A-type potassium ($\bar{g}_{\mathrm{A}}$) and slow calcium ($\bar{g}_{\mathrm{CaS}}$) channels to transition neurons from tonic spiking to bursting, and then allowed calcium-homeostatic compensation to unfold. In all illustrative simulations (Fig 2), the time constant of the calcium-homeostatic controller was reduced so that the dynamics of both neuromodulation and calcium-homeostatic compensation could be observed within a single trace, without affecting the qualitative behavior of the system.
Fig 2: Sharp neuromodulation interferes with homeostasis whereas controlled neuromodulation leads to robust and modulable neuronal function. A. Time evolution of all conductances in the STG model, displayed on a logarithmic scale, during calcium homeostasis with sharp neuromodulation (instantaneous changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line) for a degenerate population of $N=200$ neuron models. B. The corresponding mean intracellular calcium concentration over time shows the target value (red dash-dotted line) is maintained both before and after sharp neuromodulation. C. Three representative traces of neurons within the population are highlighted (top) before neuromodulation ($t_{\text{before}}$), right after ($t_{\text{after}}$) and after calcium-homeostatic compensation ($t_{\text{compensated}}$); along with the activity distributions of the population at these three time points (bottom). On the one hand, before neuromodulation, the population exhibits tonic firing, and hence the firing frequency (defined as the reciprocal of the interspike interval) is used as the activity metric. On the other hand, after neuromodulation, the population exhibits bursting, and hence the burstiness is used as the activity metric. Burstiness is defined as in [14], i.e., the product of the intraburst frequency (reciprocal of the interspike interval within a burst), the interburst frequency (reciprocal of the interburst interval), and the number of spikes per burst. A burstiness of zero means that the neuron is not producing multi-spike bursts (silence, tonic firing, or single-spike bursters). D. Same as panel A, but with controlled neuromodulation (controlled changes in $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ starting at the dashed blue line). E. Same as panel B, but with controlled neuromodulation. F. Same as panel C, but with controlled neuromodulation.
We first test the interaction of calcium homeostasis and sharp neuromodulation (Fig 2A-C). All conductance trajectories ($N=200$) are shown in Fig 2A, while Fig 2B illustrates calcium dynamics across the population. Specific time evolution curves for single neurons within the population can be found in Supporting Information, S1 Appendix. Despite initially similar firing patterns ($t_{\text{before}}$) and a similar outcome of sharp neuromodulation ($t_{\text{after}}$), calcium-homeostatic compensation produced highly variable and often unreliable outcomes in terms of modulated activity ($t_{\text{compensated}}$), while still regulating intracellular calcium (Fig 2C). Depending on the initial conductance state, neuromodulation led to new states where calcium homeostasis could either maintain functional activity, result in pathological behavior (loss of bursting, i.e., a burstiness of 0), or drive neurons into unphysiological regimes. This finding underscores a fundamental property of degeneracy: neurons with different underlying conductances can exhibit similar activity yet respond differently to perturbations, such as sharp neuromodulation. Note that these differences in channel conductances are consistent with biological measurements: different channel types have effective conductances spanning several orders of magnitude [50, 45]. Moreover, the effective conductance of a given ion channel type can vary severalfold within a single neuron type [57]. The variability in the dataset from [15] is inspired by these biological measurements and slightly exceeds the observed range to test the robustness of our analysis.
This type of interference was anticipated in [45], where the authors demonstrated that calcium-homeostatic regulation can produce diverse responses to perturbations such as ion channel deletions. Depending on the initial conditions, a deletion may either preserve function, disrupt it, or lead to pathological compensation. Our results extend this idea by showing that even physiological disturbances, such as neuromodulation, can induce similar unpredictability when coupled with calcium-homeostatic regulation.
@@
Using controlled neuromodulation on the same population dramatically changes the outcome (Fig 2D-F). Fig 2D-E illustrates the transition from tonic spiking to bursting in the same degenerate population of STG neuron models as in Fig 2A-C. Specific time evolution curves for single neurons within the population can be found in Supporting Information, S1 Appendix. Initially, all neurons exhibit tonic firing, with calcium homeostasis tuning conductances along a tonic spiking homogeneous scaling line to maintain a target calcium level ($t_{\text{before}}$ in Fig 2F). Midway through the simulation, controlled neuromodulation is applied, controlling $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$ to induce bursting ($t_{\text{after}}$ in Fig 2F). As in the case of sharp neuromodulation, this transition increases intracellular calcium levels, leading to an adaptive shift in conductance values. Calcium-homeostatic regulation then readjusts overall conductance magnitudes, but here the interaction between calcium homeostasis and controlled neuromodulation ensures that the newly established neuromodulated state is maintained. Ultimately, the population reaches a stable configuration where both the neuromodulated function and intracellular calcium homeostasis are fulfilled ($t_{\text{compensated}}$ in Fig 2F).
The differences between sharp and controlled neuromodulation can also be highlighted through neuromodulator washout, i.e., removal of the neuromodulatory input. At both $t_{\text{after}}$ and $t_{\text{compensated}}$, washout recovers tonic spiking across the population under both sharp and controlled neuromodulation, though for fundamentally different reasons and with different levels of fidelity. Under sharp neuromodulation, calcium-homeostatic compensation continuously drives conductances along the native tonic spiking scaling direction, so that tonic spiking is the only remaining stable activity pattern. Recovery of spiking upon washout is therefore expected, as the system has no other stable attractor. However, when washout occurs at $t_{\text{compensated}}$, the conductance distribution of the population does not exactly match its pre-neuromodulation state: because calcium homeostasis has been compensating for the sharp neuromodulatory perturbation throughout, the resulting conductance profiles are subtly altered, leading to a slightly different distribution of firing frequencies across the population. Although the qualitative behavior (tonic spiking) is preserved, this quantitative drift implies that repeated cycles of sharp neuromodulation and washout could progressively erode the fidelity of neuronal function, as each cycle introduces small but cumulative changes in the conductance landscape. In contrast, under controlled neuromodulation, the system can maintain any target activity through activity-dependent feedback. Upon washout, the neuromodulatory reference is removed, and the system returns to tonic spiking because the calcium-homeostatic controller regains sole authority over conductance tuning. Because the controlled neuromodulation mechanism continuously coordinates with calcium homeostasis rather than competing with it, the pre-neuromodulation conductance distribution is faithfully recovered upon washout.
@@
Controlled neuromodulation tunes the direction of calcium homeostasis-driven changes in conductance values to ensure stable neuromodulated function during calcium-homeostatic regulation
@@
To study the mechanisms underlying the constructive interaction between controlled neuromodulation and calcium homeostasis, we plot the conductance trajectories of the simulations from Fig 2 in the plane of modulated conductances, $\bar{g}_{\mathrm{CaS}}$ and $\bar{g}_{\mathrm{A}}$, during tonic spiking calcium-homeostatic compensation for both sharp and controlled neuromodulation (Fig 3A). Initially, as calcium levels remain below the threshold, calcium homeostasis increases the modulated conductances along a homogeneous scaling direction characteristic of tonic spiking. This direction corresponds to a line intersecting the origin, with variability due to conductance degeneracy within the population, and its slope is determined by the relative time constants of mRNA transcription in the calcium homeostasis model, with steady-state analysis predicting $\bar{g}_{i}/\bar{g}_{j}\approx\tau_{m_{j}}/\tau_{m_{i}}$ [45]. When neuromodulation is applied, all conductances are adjusted to reshape their ratios and induce bursting (Fig 3B). Importantly, neuromodulatory actions follow the same direction across neurons but vary in magnitude depending on individual neuronal states [15]. This modulation leads to a sharp increase in intracellular calcium levels. After neuromodulation is applied, the long-term outcome depends on the type of neuromodulation used (Fig 3C-D).
@@
Fig 3: Controlled neuromodulation tunes the direction of homeostasis-driven changes to ensure stable neuromodulated function. A. Trajectories of the STG population from Fig 2 in the modulated conductance space during homogeneous scaling in tonic spiking only. As expected from calcium homeostasis, all neurons move in the same direction. White circles represent the initial conditions at the start of the step, and the green triangle marks the final conditions. B. Same as panel A, but during the brief period when neuromodulation is active. All trajectories are parallel to one another. C. Same as panels A and B, but during the calcium homeostasis compensation phase with sharp neuromodulation, resulting in non-robust outcomes. D. Same as panel C, but with controlled neuromodulation, resulting in function-preserving robust outcomes.